OpenStudy (itsbribro):

What is the equation of a line that passes through (4, 3) and has a slope of 2?

OpenStudy (legomyego180):

SLOPE INTERCEPT FORM Y=MX+B

OpenStudy (itsbribro):

A. y=2x-11 B. y=2x-5 C. y=2x-1 D. y=2x-7

OpenStudy (itsbribro):

so can you fit the numbers in for me I got confused

OpenStudy (legomyego180):

you can always just try plugging the numbers in too

OpenStudy (legomyego180):

x=4 y=3 plug in x=4 for each answer and the answer that gives you e as your y value is the correct one!

OpenStudy (itsbribro):

3=2(4)+b?

OpenStudy (itsbribro):

omg... i'm confused

OpenStudy (legomyego180):

3 not e. e is a number too but you'll learn about that later.

OpenStudy (itsbribro):

y=2x+3?

OpenStudy (itsbribro):

@princeharryyy please help..

OpenStudy (mrnood):

@itsbribro just take a moment and slow down a bit

OpenStudy (legomyego180):

\[(4, 3)\] x=4, y=3 A. \(2(4)-11=-3\) B. \(2(4)-5=3\) C. \(2(4)-1=7\) D. \(2(4)-7=1\)

OpenStudy (owlcoffee):

\[(y-y_o)=m(x-x_o)\] Just take the point, plug the data in and put it into the y=mx+c form: \[(y-3)=2(x-2)\]

OpenStudy (mrnood):

The equation you want LOOKS like y = mx + b where m is the slope so can you write that equation (leaving x and y) with YOUR value of m?

OpenStudy (itsbribro):

thanks guys..

OpenStudy (legomyego180):

>.< We can teach you but we cant learn for you

OpenStudy (princeharryyy):

Mr nood way is easy

OpenStudy (princeharryyy):

Maybe someday @legomyego180 we might.

OpenStudy (itsbribro):

i know i was just asking you how i would plug the numbers in. i was getting confused the way you were explaining @legomyego180

OpenStudy (itsbribro):

thank you all for your help explaining this to my stupid brain.. xoxo

OpenStudy (legomyego180):

Mr. Noods way is the correct way. My way was kind of the easy way out, if you will. you have (x,y) where and and y are numbers lets say you are given (1,4) with a slope of 2 *Your slope is always going to be the number in front of your x (i.e. 7x, 10x, 1x) A. y=2x+1 B. y=2x-4 C. y=2x+2 D. y=2x+8 From your coordinates (1,4) you see that x=1. So in each of your answers plug in 1 for the x value and see what you get out. Ill do it below A. y=2(1)+1 =3 B. y=2(1)-4 = -2 C. y=2(1)+2 =4 D. y=2(1)+8 =10 Now looking back at your coordinated (1,4) you see that your y-value has to equal 4 The only answer that equaled 4 after you plugged in 1 for x was C So for this example question C would be your answer. Its not something necessary to know, just a good way to check your work, although it doesnt work 100% of the time, which is why my noods way is better